The sum of my digits is 13. My hundreds place is 3 less than 5. My ones place is 4 more than 3. My thousands place is even. What 4-digit numbers can I be?
Found 2 solutions by ReadingBoosters, stanbon:
Answer by ReadingBoosters(3246)

t = thousands
x = hundreds
y = tens
z = ones
t+x+y+z=13
x=5-3 = 2
z=3+4 = 7
t + y = 13 - 2 - 7
t + y = 4
t is even
...
2,227
4,207
247
...............
Delighted to help
-Reading Boosters
Wanting for others what we want for ourselves.

Answer by stanbon(75887)

The sum of my digits is 13. My hundreds place is 3 less than 5. My ones place is 4 more than 3. My thousands place is even. What 4-digit numbers can I be?
-----
Let the number be abcd
Equations:
a + b + c + d = 13
b = 2
d = 7
a = 2x
--------
Substitute:
2x + 2 + c + 7 = 13
2x + c = 4
--
Since 2x is even, x = 1 and c = 2
---
Ans: 2227
===============
Cheers,
Stan H.